# On variants of Chebyshev’s conjecture

### v2

@@ 4.3. Proof of Theorem 5
It suffices to show (1.20) with $x$ replaced by $x/e^{2}$. Using the decomposition
@@
and (2.1), it can be obtained by an argument similar to that of Theorem 4. $\Box$
@@ 5.1. Proof of Theorem 6
we take the sum of (4.4) over all $\chi$ modulo $q$. Then,
$\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)=\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}.$ (5.1)
The Dirichlet theorem on arithmetic progressions [15, Corollary 11.17] tells us that the leading term in the asymptotic formulas for both (5.1) and
@@
$f_{q}(x)$ $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)-\frac{4\sqrt{x}}{\varphi(q)}$ (5.2)
@@
$F_{q}(x)$ $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)-\frac{2}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}$ (5.3)
@@
Taking the sum $\varphi(q)^{-1}\sum_{\chi\,{\rm mod}\,q}$ in (4.5),
$\int_{1}^{\infty}$ $(-f_{q}(x))x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{\varphi(q)}\left(\frac{1}{(s-1/2)^{2}}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}(s,\chi)+\frac{4}{s-1}\right)$ (5.4)
for $\Re(s)>1$. On the other hand, by (3.6), (4.5), and the way of their proofs, we obtain
$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\right)x^{-s+1/2}\,\frac{dx}{x}=\frac{2(1-s)}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s,\chi)$ (5.5)
for $\Re(s)>1$. Taking the sum $\varphi(q)^{-1}\sum_{\chi\,{\rm mod}\,q}$ in (5.5),
$\int_{1}^{\infty}(-F_{q}(x))x^{-s+1/2}\,\frac{dx}{x}=\frac{2(1-s)}{(s-1/2)^{2}}\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}(s,\chi).$ (5.6)
Then, following an argument similar to the proof of Theorems 2 and 8, and using (5.4) and (5.6), we obtain the first claim of Theorem 6.
We now prove the second and third claims of Theorem 6. Suppose that $L(1/2,\chi)\not=0$ for any Dirichlet character $\chi$ modulo $q$. In an argument similar to the proof of (3.8), in the case $\chi=\chi_{0}$, we include the zeros of $\prod_{p|q}(1-p^{-s})$ in both sums for $\rho$ on the right-hand side of (3.8); in the case $\chi\neq\chi_{0}$, we use (4.8) and (2.13) instead of (3.1) and (2.6), respectively, and include the zeros of $\prod_{p|q}(1-\chi^{\ast}(p)p^{-s})$ in both sums for $\rho$ on the right-hand side, to obtain
$\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}$ $=-\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)\log x+2\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}$ (5.7)
@@
Here, the two sums $\sum_{\rho_{\chi}}$ on the right-hand side are taken over all nontrivial zeros of $L(s,\chi^{\ast})$ and all zeros of the finite product $L(s,\chi)/L(s,\chi^{\ast})=\prod_{p\mid q}(1-\chi^{\ast}(p)p^{-s})$, $\kappa^{\ast}=\kappa(\chi^{\ast})$, and the last term arises from the difference between $\sideset{}{{}^{\prime}}{\sum}_{n\leq x}$ and $\sum_{n\leq x}$. Taking the sum $\sum_{\chi\,{\rm mod}\,q}$ in (4.8) and using (3.1) for $\chi=\chi_{0}$, we obtain
$\varphi(q)f_{q}(x)$ $=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)-\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}-1/2}}{(\rho_{\chi^{\ast}}-1/2)^{2}}+O(1),$ (5.8)
where the sums $\sum_{\rho_{\chi^{\ast}}}$ on the right-hand side are taken over all nontrivial zeros of $L(s,\chi^{\ast})$ and the implied constant depends only on $q$. Similarly, taking the sum $\sum_{\chi\,{\rm mod}\,q}$ in (5.7), we obtain
$\varphi(q)F_{q}(x)$ $=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$ (5.9)
@@
where the sums $\sum_{\rho_{\chi}^{\ast}}$ and the implied constant have the same meaning as in (5.8). In (5.9), the contribution from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ is included in the $O(1)$ term, and it is justified as follows.
For a Dirichlet character $\chi$ modulo $q$, we denote by $\mathcal{Z}_{1}(\chi)$ the set of all nontrivial zeros of $L(s,\chi^{\ast})$ counted with multiplicity and by $\mathcal{Z}_{2}(\chi)$ the set of all zeros of $L(s,\chi)/L(s,\chi^{\ast})$. Then, the second sum on the right-hand side of (2.13) for $s=1/2$ is written as
@@
is convergent and bounded, as in the case of the Fourier series of the sawtooth wave. Therefore, the contribution from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ on the right-hand side of (5.7), which originates from (2.13), is bounded. Moreover, the contributions from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ arising from (3.1) and (4.8) are also bounded, since the sums over zeros in (3.1) and (4.8) converge absolutely.
Assuming that the GRH holds for $L(s,\chi)$ for all $\chi$ modulo $q$, the second sum on the right-hand side of (5.8) is bounded. Additionally, if we assume (1.25), the last sum over $\chi$ mod $q$ on the right-hand side of (5.9) is $o(\log x)$. Hence, (1.23) and (1.24) hold.
Conversely, assuming (1.23) (resp. (1.24)), it follows from (5.2) (resp. (5.3)) that $f_{q}(x)\ll\log x$ (resp. $F_{q}(x)\ll\log x$) as $x\to\infty$. Therefore, the integral on the left-hand side of (5.4) (resp. (5.6)) converges uniformly on any compact subset of the right half-plane $\Re(s)>1/2$, and hence defines an analytic function there. Thus, the GRH for $L(s,\chi)$ for all $\chi$ modulo $q$ follows from an argument similar to that in the proof of Theorem 8. Moreover, assuming (1.24), it follows from (5.3) that
@@
Comparing this with (5.9), we find that the sum over $\chi$ mod $q$ on the right-hand side of (5.9) must be $o(\log x)$. However, the latter part of this sum is absolutely convergent and bounded, and thus (1.25) follows.
Finally, we prove the fourth and fifth claims of Theorem 6. By using (2.14) in place of (2.13) and (4.10) in place of (4.8), and employing the resulting modified versions of (5.8) and (5.9), one can show that both (1.26) and (1.27) are equivalent to the GRH for $L(s,\chi)$ for all Dirichlet characters $\chi$ modulo $q$, by an argument similar to that used in the proofs of the second and third claims. Furthermore, by appropriately modifying (5.9) using (2.14) and (4.10), one can also show that (1.27) implies (1.28) by an argument similar to that of the proof of the third claim. We omit the details, as there is no essential difference. $\Box$
@@ 5.2. Proof of Theorem 7
First, we prove (1.29). After multiplying $f_{q}(x)$ in (5.2) by $\varphi(q)$, summing over $3\leq q\leq Q$, and then changing the order of summation, we obtain
$\sum_{3\leq q\leq Q}\varphi(q)f_{q}(x)$ $=\sum_{3\leq q\leq Q}\left(\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\sqrt{x}\right)$ (5.10)
@@
where we understand that the inner sum is zero if it is an empty sum. For any $n\leq x$, all the divisors of n1n-1 appear in the range $1\leq q\leq Q$ by the assumption $Q\geq x$. For a prime number $p$, 22 divides $p^{k}-1$ if and only if $p$ is odd. Therefore, using $\sum_{d|m}\varphi(d)=m$ on the right-hand side of (5.10),
@@
Using [21, (2.3)] and the prime number theorem (2.2) in (3.1),
@@
By applying partial summation to the first term on the right-hand side using the prime number theorem (2.2),
@@
Hence, we obtain (1.29).
Next, we prove (1.30). After multiplying $F_{q}(x)$ in (5.3) by $\varphi(q)$, summing over $3\leq q\leq Q$, and then changing the order of summation as in (5.10), we obtain
$\sum_{3\leq q\leq Q}\varphi(q)F_{q}(x)=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\sum_{{3\leq q\leq Q}\atop{q\mid n-1}}\varphi(q).$ (5.11)
@@
Using [21, (2.3)] and the prime number theorem (2.2) in (3.8),
@@
By applying partial summation to the first term on the right-hand side using the prime number theorem (2.2),
@@
Hence, we obtain (1.30) by substituting $xe^{2}$ for $x$. $\Box$
If we take the sum $\sum_{3\leq q\leq Q}$ in (5.3), we obtain
@@
This also shows that (1.22) has a constant sign on average, however, the sum used in the proof of Theorem 7 is probably easier to handle.

### v3

@@ 4.3. Proof of Theorem 5
It suffices to show (20) with $x$ replaced by $x/e^{2}$. Using the decomposition
@@
and (31), it can be obtained by an argument similar to that of Theorem 4. $\Box$
@@ 5.1. Proof of Theorem 6
we take the sum of (58) over all $\chi$ modulo $q$. Then,
$\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)=\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}.$ (72)
The Dirichlet theorem on arithmetic progressions [15, Corollary 11.17] tells us that the leading term in the asymptotic formulas for both (72) and
@@
$f_{q}(x)$ $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)-\frac{4\sqrt{x}}{\varphi(q)}$ (73)
@@
$F_{q}(x)$ $=\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}f_{\chi}(x)-\frac{2}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}$ (74)
@@
Taking the sum $\varphi(q)^{-1}\sum_{\chi\,{\rm mod}\,q}$ in (59),
$\int_{1}^{\infty}$ $(-f_{q}(x))x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{\varphi(q)}\left(\frac{1}{(s-1/2)^{2}}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}(s,\chi)+\frac{4}{s-1}\right)$ (75)
for $\Re(s)>1$. On the other hand, by (51), (59), and the way of their proofs, we obtain
$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\right)x^{-s+1/2}\,\frac{dx}{x}=\frac{2(1-s)}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s,\chi)$ (76)
for $\Re(s)>1$. Taking the sum $\varphi(q)^{-1}\sum_{\chi\,{\rm mod}\,q}$ in (76),
$\int_{1}^{\infty}(-F_{q}(x))x^{-s+1/2}\,\frac{dx}{x}=\frac{2(1-s)}{(s-1/2)^{2}}\frac{1}{\varphi(q)}\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}(s,\chi).$ (77)
Then, following an argument similar to the proof of Theorems 2 and 8, and using (75) and (77), we obtain the first claim of Theorem 6.
We now prove the second and third claims of Theorem 6. Suppose that $L(1/2,\chi)\not=0$ for any Dirichlet character $\chi$ modulo $q$. In an argument similar to the proof of (53), in the case $\chi=\chi_{0}$, we include the zeros of $\prod_{p|q}(1-p^{-s})$ in both sums for $\rho$ on the right-hand side of (53); in the case $\chi\neq\chi_{0}$, we use (62) and (43) instead of (46) and (36), respectively, and include the zeros of $\prod_{p|q}(1-\chi^{\ast}(p)p^{-s})$ in both sums for $\rho$ on the right-hand side, to obtain
$\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}$ $=-\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)\log x+2\sum_{\rho_{\chi}}\frac{x^{\rho_{\chi}-1/2}}{\rho_{\chi}-1/2}$ (78)
@@
Here, the two sums $\sum_{\rho_{\chi}}$ on the right-hand side are taken over all nontrivial zeros of $L(s,\chi^{\ast})$ and all zeros of the finite product $L(s,\chi)/L(s,\chi^{\ast})=\prod_{p\mid q}(1-\chi^{\ast}(p)p^{-s})$, $\kappa^{\ast}=\kappa(\chi^{\ast})$, and the last term arises from the difference between $\sideset{}{{}^{\prime}}{\sum}_{n\leq x}$ and $\sum_{n\leq x}$. Taking the sum $\sum_{\chi\,{\rm mod}\,q}$ in (62) and using (46) for $\chi=\chi_{0}$, we obtain
$\varphi(q)f_{q}(x)$ $=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)-\sum_{\chi\,{\rm mod}\,q}\sum_{\rho_{\chi^{\ast}}}\frac{x^{\rho_{\chi^{\ast}}-1/2}}{(\rho_{\chi^{\ast}}-1/2)^{2}}+O(1),$ (79)
where the sums $\sum_{\rho_{\chi^{\ast}}}$ on the right-hand side are taken over all nontrivial zeros of $L(s,\chi^{\ast})$ and the implied constant depends only on $q$. Similarly, taking the sum $\sum_{\chi\,{\rm mod}\,q}$ in (78), we obtain
$\varphi(q)F_{q}(x)$ $=-\log x\sum_{\chi\,{\rm mod}\,q}\frac{L^{\prime}}{L}\left(\frac{1}{2},\chi\right)$ (80)
@@
where the sums $\sum_{\rho_{\chi}^{\ast}}$ and the implied constant have the same meaning as in (79). In (80), the contribution from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ is included in the $O(1)$ term, and it is justified as follows.
For a Dirichlet character $\chi$ modulo $q$, we denote by $\mathcal{Z}_{1}(\chi)$ the set of all nontrivial zeros of $L(s,\chi^{\ast})$ counted with multiplicity and by $\mathcal{Z}_{2}(\chi)$ the set of all zeros of $L(s,\chi)/L(s,\chi^{\ast})$. Then, the second sum on the right-hand side of (43) for $s=1/2$ is written as
@@
is convergent and bounded, as in the case of the Fourier series of the sawtooth wave. Therefore, the contribution from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ on the right-hand side of (78), which originates from (43), is bounded. Moreover, the contributions from the zeros of $L(s,\chi)/L(s,\chi^{\ast})$ arising from (46) and (62) are also bounded, since the sums over zeros in (46) and (62) converge absolutely.
Assuming that the GRH holds for $L(s,\chi)$ for all $\chi$ modulo $q$, the second sum on the right-hand side of (79) is bounded. Additionally, if we assume (25), the last sum over $\chi$ mod $q$ on the right-hand side of (80) is $o(\log x)$. Hence, (23) and (24) hold.
Conversely, assuming (23) (resp. (24)), it follows from (73) (resp. (74)) that $f_{q}(x)\ll\log x$ (resp. $F_{q}(x)\ll\log x$) as $x\to\infty$. Therefore, the integral on the left-hand side of (75) (resp. (77)) converges uniformly on any compact subset of the right half-plane $\Re(s)>1/2$, and hence defines an analytic function there. Thus, the GRH for $L(s,\chi)$ for all $\chi$ modulo $q$ follows from an argument similar to that in the proof of Theorem 8. Moreover, assuming (24), it follows from (74) that
@@
Comparing this with (80), we find that the sum over $\chi$ mod $q$ on the right-hand side of (80) must be $o(\log x)$. However, the latter part of this sum is absolutely convergent and bounded, and thus (25) follows.
Finally, we prove the fourth and fifth claims of Theorem 6. By using (44) in place of (43) and (64) in place of (62), and employing the resulting modified versions of (79) and (80), one can show that both (26) and (27) are equivalent to the GRH for $L(s,\chi)$ for all Dirichlet characters $\chi$ modulo $q$, by an argument similar to that used in the proofs of the second and third claims. Furthermore, by appropriately modifying (80) using (44) and (64), one can also show that (27) implies (28) by an argument similar to that of the proof of the third claim. We omit the details, as there is no essential difference. $\Box$
@@ 5.2. Proof of Theorem 7
First, we prove (29). After multiplying $f_{q}(x)$ in (73) by $\varphi(q)$, summing over $3\leq q\leq Q$, and then changing the order of summation, we obtain
$\sum_{3\leq q\leq Q}\varphi(q)f_{q}(x)$ $=\sum_{3\leq q\leq Q}\left(\varphi(q)\sum_{{n\leq x}\atop{n\equiv 1\,{\rm mod}\,q}}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-4\sqrt{x}\right)$ (81)
@@
where we understand that the inner sum is zero if it is an empty sum. For any $n\leq x$, all the divisors of n1n-1 appear in the range $1\leq q\leq Q$ by the assumption $Q\geq x$. For a prime number $p$, 22 divides $p^{k}-1$ if and only if $p$ is odd. Therefore, using $\sum_{d|m}\varphi(d)=m$ on the right-hand side of (81),
@@
Using [21, (2.3)] and the prime number theorem (32) in (46),
@@
By applying partial summation to the first term on the right-hand side using the prime number theorem (32),
@@
Hence, we obtain (29).
Next, we prove (30). After multiplying $F_{q}(x)$ in (74) by $\varphi(q)$, summing over $3\leq q\leq Q$, and then changing the order of summation as in (81), we obtain
$\sum_{3\leq q\leq Q}\varphi(q)F_{q}(x)=\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\sum_{{3\leq q\leq Q}\atop{q\mid n-1}}\varphi(q).$ (82)
@@
Using [21, (2.3)] and the prime number theorem (32) in (53),
@@
By applying partial summation to the first term on the right-hand side using the prime number theorem (32),
@@
Hence, we obtain (30) by substituting $xe^{2}$ for $x$. $\Box$
If we take the sum $\sum_{3\leq q\leq Q}$ in (74), we obtain
@@
This also shows that (22) has a constant sign on average, however, the sum used in the proof of Theorem 7 is probably easier to handle.